OpenStudy (anonymous):

Solve each equation. Show all your work. Round your answers to four decimal places. 4.) 74x = 415 17.) log7 (x +1) + log7(x – 5) = 1 I'm only stuck on these two problems, help please!!

OpenStudy (anonymous):

4.) 7^4x = 415, sorry.

OpenStudy (owlcoffee):

Okay, let's look at : \[7^{4x}=415\] Now, what we awant to do here, is create a exponent with the same base or but let's apply logarithms on both sides of the equality, ending with: \[\log_{7} 7^{4x}=\log_{7} 415\] Now, we have a property of the logarithms that I won't prove, but I'll show it to you: \[\log_{}A ^{n} =n \log_{}A \] The property states that if I have a logarithm number with an exponent, it's the same thing as making it a product with the exponent. So let's apply that: \[\log_{7} 7^{4x}=4x \log_{7} 7\] But Logarithm with base 7 of 7 is equal to 1, meaning that: \[4x \log_{7} 7=4x\] \[\log_{7} 7=1\] Let's replace what we found out on the original equation: \[4x=\log_{7} 415\] Now I want to leave the x alone on the left side of the equality, so I'll divide both sides by 4: \[\frac{ 4x }{ 4 }=\frac{ \log_{7}415 }{ 4 }\] now the 4's on the left side will simplify and the right side is equal to a constant. \[x=\frac{ \log_{7} 415 }{ 4 }\]

OpenStudy (anonymous):

so do i have to divide log7 415 and 4 to get x?

OpenStudy (owlcoffee):

That's correct.

OpenStudy (anonymous):

and what is it that log7 equals ?

OpenStudy (jdoe0001):

\(\bf 7^{4x} = 415\implies log_7(7^{4x})=log_7415\implies 4x=log_7415\\ \quad \\ \textit{change of base rule of }{\color{blue}{ log_ab=\cfrac{logcb}{log_ca}}}\\ 4x=log_7415\implies x=\cfrac{log_7415}{4}\implies \large x=\cfrac{ {\color{blue}{ \frac{log_{10}415}{log_{10}7}}} }{4}\)

OpenStudy (jdoe0001):

ahemm \(\bf 7^{4x} = 415\qquad \textit{log cancellation rule of } {\color{blue}{ log_aa^x=x}}\) btw =)

OpenStudy (owlcoffee):

There's no need for a change of base, Log7 (415) over 4, is already a constant. We have no variables in game.

OpenStudy (jdoe0001):

well.... true

OpenStudy (owlcoffee):

But I believe you wanted to make it more, calculator friendly. Thank you ~~

OpenStudy (anonymous):

does it equal 7 because it cancels out? 7(415) / 4 = 726.25 x = 726.25? im horrible with math.

OpenStudy (owlcoffee):

\[x=\frac{ \log_{7}415 }{ 4 }\] \[\log_{7} 415= 3.097922\] \[=> x=\frac{ 3.097922 }{ 4 }\] x=0.7744

OpenStudy (jdoe0001):

as far as 17) \(\bf recall\qquad {\color{blue}{ log(a)+log(b)\implies log(a\cdot b)}}\qquad thus\\ \quad \\ log_7(x +1) + log_7(x - 5) = 1\implies log_7[(x +1)(x - 5)]=1\\ \quad \\ \textit{log cancellation rule of }\Large {\color{red}{ a}}^{log_{\color{red}{ a}}x}=x\qquad thus\\ \quad \\ log_7[(x +1)(x - 5)]=1\implies\Large {\color{red}{ 7}}^{log_{\color{red}{ 7}}[(x +1)(x - 5)]}={\color{red}{ 7}}^1 \\ \quad \\ (x +1)(x - 5)=7\)

OpenStudy (jdoe0001):

as you'd end up with a quadratic, set it to 0, factor it to get the values for "x"

OpenStudy (anonymous):

ohhhhh, okay. thank you!! and would you be able to help with the next one? (:

OpenStudy (jdoe0001):

I'm about to dash, will be here tomorrow though :)